Speech delivered at CADE13 Business Meeting

David A. Plaisted

Today we are deciding between the current system, in which none of the trustees are elected by the membership, the trustees’ proposal, in which two-thirds or less of them are elected, and my proposal, in which all of them are elected by the membership.

For the sake of simplicity, I will assume that CADE will meet once a year. In my proposed system, the trustees would have three year terms, three elected each year, making nine trustees in all. The trustees could also appoint other officers, such as a secretary and a treasurer, for an unlimited number of consecutive terms, but these officers would not automatically be trustees. As in the current system, the trustees would choose the CADE program chairs and local arrangements chairs, but these chairs would not become trustees unless they were elected. The trustee election would be by a simple two-round system, in which the second round could be omitted by a vote of the membership. Each member would cast three votes for the three vacant trustee positions. The proposed system also would permit special elections, in which all members of the AAR could vote. These special elections would occur between CADE conferences, and would help to reduce the effects of locality caused by the possibly small percentage of AAR members at each CADE conference. Motions passed at business meetings would be binding on the trustees, but could be reversed in a special election, which can reach the entire constituency. Our general philosophy has been to change the current system as little as possible and retain its good features, while making it fully democratic. We have made every effort to accommodate the concerns of the trustees in the development of this proposal.

The trustees’ proposal has somewhere between 3 and 6 non-elected trustees as well as six elected trustees, reducing the democratic portion. It permits the trustees to cancel motions passed at business meetings and special elections, if they choose. This is especially disturbing for special elections, which reach the whole constituency via e mail. It has too high a proportion (66 percent) for changing the bylaws. My proposal requires only a simple majority to revise the bylaws. It takes a long time to understand the implications of the bylaws, and we should be flexible at the beginning. The means proposed for conducting elections in the trustees’ proposal is a lot of work and not necessarily any better than my two-round system. Their voting system also means that the results of the election will not be known until later. Also, with two trustees elected each year, it is likely that the trustee positions would be filled with program chairs serving their first and second terms, reducing the chance for the membership to elect an additional trustee.

I don’t believe we need to be so afraid of what attendees at one CADE conference will do as to make their motions non binding on the trustees. I think that people will be reasonable in how they vote, and if mistakes are made, they can be corrected in a special election.

The trustees criticized one version of my proposal because it excluded some people from the slate. This feature has been removed.

I would like to pay particular attention to the question of whether upcoming program chairs should be voting trustees even without being elected. This serves no real purpose, and diminishes the democratic aspect of the system. In my plan, the upcoming program chairs could participate in all relevant trustee discussions and have an influence even without a vote. I believe that many program chairs would be willing to give up their vote among the trustees in order to insure a fully democratic system. Furthermore, in all three plans, program chairs and the secretary/treasurer will be voting members of the conference committee that has charge of their CADE conference.

Finally, I believe that it is a mistake for upcoming program chairs and the secretary/treasurer to have a vote in the choice of future program chairs.

A democratic system doesn’t mean 6 elected and 3 or more non-elected positions; it means that all of the positions are elected. I am opposed to any attempt to reduce the democratic portion of the system. It doesn’t matter how well or how badly people are doing their jobs or how successful or unsuccessful things are. Democracy should not be subject to negotiation.

In the current system, there is no regular procedure for evaluating program chairs. Whether good or bad, program chairs become trustees, and help to choose future program chairs. There is a mechanism for removing program chairs, but it would probably be used only very rarely due to the embarrasment it would cause, and even then, the current trustees would do the evaluation.

Some other conferences have non-democratic systems, but their members might also vote to make them democratic if they were given a chance. In fact, I would recommend that all conferences should be made democratic.

I don’t claim that a democratic system will solve all of the problems of CADE, whatever they may be. But it will be an improvement, and with very little cost attached to it.

I believe we are at a turning point in the history of CADE. The current trustees will pass off the scene, but the organizational structure of CADE will last a long time. If we do not make the change to a democratic system now, the current system will become more and more entrenched and harder to change. It’s unlikely that anyone else will go to the considerable effort required to change to a democratic system. If we wait until a serious problem develops, then we will have to deal with the problem and also try to change to a democratic system at the same time, which could be very difficult to do. Now when things are running relatively smoothly is the best time to make the change. The issue is not whether you think the trustees are doing a good job - if you think that, you are free to vote for them under a democratic system. The issue is not the personnel, but the system. I encourage you to vote in favor of the fully democratic bylaws.